# Femtoscopy analysis of ultra-soft pion trap at LHC energies

### v1

@@ I INTRODUCTION
In experimental studies, one clearly observes certain types of universal scaling behaviour for the measured femtoscopy scales. For a given colliding system and collision energy a linear scaling of the femtoscopic radii is universally observed versus cube root of the final state mean particle multiplicity $\langle dN_{ch}/d\eta\rangle^{1/3}$ [16, 17, 18, 19, 20, 21, 22]. A similar scaling across collision energies and colliding systems is only approximate [23, 24]. The radii versus the pair transverse momentum exhibit a power-law-like scaling in pair transverse mass $m_{T}$ ($m_{T}$ = $\sqrt{k_{T}^{2}+m^{2}}$, where $k_{T}$ is pair mean transverse momentum, $k_{T}=|\textbf{p}_{T1}+\textbf{p}_{T2}|$/2, and $m$ is particle mass) [25, 26, 27, 28, 29, 16, 17, 30]. In heavy-ion collisions both these scalings are predicted by hydrodynamic models [31]. Specifically, the $m_{T}$ scaling is explained as a direct consequence and one of the main signatures of the collective radial flow of the system.
@@
Thus, in this paper, we focus on studying the pion femtoscopic radii dependence on pair $m_{T}$, including the region of very low transverse masses, in Pb-Pb collisions at the LHC energy $\sqrt{s_{NN}}$ = 5.02 TeV simulated within the integrated HydroKinetic Model (iHKM) [31, 38, 39] and the hydrodynamics model [40] coupled to the statistical hadronisation code THERMINATOR2 (LHYQUID + THERMINATOR2) [41]– (LQTH).
@@ II MODELS’ DESCRIPTION
The LQTH approach assumes the ideal (3+1)D hydrodynamics evolution of the created fireball (however, bulk and shear viscosity can be implemented as well) until the single (both chemical and kinetic) freeze-out at certain temperature ($T_{p}\approx 140$ MeV), where statistical hadronization takes place, and the system’s chemical composition and momentum spectra do not change anymore, except for the resonance decays. As well as the iHKM, for the current study the model is properly calibrated to describe the $\sqrt{s_{NN}}=5.02$ TeV Pb-Pb collisions at the LHC.
@@ III CONSTRUCTION OF THE CORRELATION FUNCTION
In experimental analysis, the procedure of obtaining the correlation function consists of making pairs of particles coming from the same event to build the correlated distribution (numerator), and for the background distribution (denominator) pairs are created using particles coming from different events (meaning that they retain all aspects of experimental acceptance, while not being correlated due to the mutual interaction).
@@ IV PION FEMTOSCOPY SCALES
This section presents the results on pion femto-radii obtained as a result of correlation function fitting for identical pions in 1D and 3D studies using the two models (iHKM and LQTH). The correlation functions were calculated separately for the six centrality classes ($0-5\%,5-10\%,10-20\%,20-30\%,30-40\%$ and $40-50\%$) and the ten $k_{T}$ ranges $(0.00-0.05,0.05-0.10,0.10-0.15,0.15-0.20,0.20-0.25,0.25-0.30,0.30-0.35,0.35-0.45,0.45-0.60,0.60-0.80)$ GeV/$c$.
@@
Figure 4: First-order derivative of the charged identical pions femtoscopic radii $m_{T}$ dependency, $dR/dm_{T}$, in the HKM (left) and the LQTH (right) models, for the three collision centralities (from top to bottom: $0-5\%,10-20\%,30-40\%$). Each panel contains derivatives of the three radii from the 3D study ($R_{\mathrm{out}}$ — blue, $R_{\mathrm{side}}$ — red, $R_{\mathrm{long}}$ — orange) and $R_{\mathrm{inv}}$ from the 1D study (green).
@@
To better investigate the character of the radii $m_{T}$ dependencies, they were differentiated, and the obtained derivatives are presented in Fig. 4 (for the three centrality classes). The derivative has been extracted in the range ±$\pm 0.07$ GeV/$c$ around each point in the figure. Most of the functions show similar behaviour, with a clear change in the derivative value at certain momenta. The iHKM model shows a constant derivative at low momentum (especially for the 1D study), then a systematic rise and again a higher constant value at large momenta. Flat parts of the curves correspond to constant slopes of the radii decrease approximated by a linear function. The rising/falling parts indicate a change in the radii dependency slope. The low-$m_{T}$ flat part ends around 0.25 GeV/$c$. The LQTH model, in turn, shows for the 1D and the 3D transverse radii derivatives a constant or decreasing behaviour up to 0.3 GeV/$c$, then the dependence starts rising to end up with a constant value at large momentum. The initial decrease reflects the change of a nearly constant $R(m_{T})$ dependence to a to decreasing one, while the next rising $dR/dm_{T}$ part is connected with more gentle radii change at growing $m_{T}$. The $long$ direction, by contrast, shows growing behaviour ending up with a plateau without any clear flat beginning.
@@
Note, that we see the mentioned effects of relative flattening of the transverse radii behaviour in the ultra-soft momentum region in “hydro plus hadronic cascade” models, where the dynamics is built in quasi-classical approximation. The Bose-Einstein quantum statistics effect is taken into account in the very final stage only, but even this way, using the femtoscopy method, one can see specific (delayed freeze-out) behaviour of emission in the ultra-low-momentum region (see also Fig. 1 from [51]). In particular, as we already mentioned, the femtoscopy analysis brings us the signal about a long-lived ultra-soft pion trap formed in the central part of the system created in ultra-relativistic heavy ion collision.
The dependencies of the radii on $m_{T}$, together with the power-law fits, are presented for the three centrality classes in Fig. 5. The fitting was performed outside the plateau region observed in Fig. 4 (above 0.25 GeV/$c$). As one can see from this figure, the power-law function in the form $am_{T}^{-b}$ works well, especially for all three $R_{i}$ in the 3D case in the iHKM model and for the $long$ direction of the LQTH model. As for the transverse radii from the LQTH, the first 2-3 corresponding points are mostly below expectations. In both models, the behaviour most different from the power-law in the low-$m_{T}$ region is observed for the 1D $R_{\mathrm{inv}}$ points, which lie, up to 0.25 GeV/$c$, below the extrapolated line.
@@
Finally, we analyzed the femto-radii dependence on pair velocity, see Fig. 7 for different centrality intervals and for both used model approaches. In all cases, we notice a slow radii falling at low velocities. All the distributions were fitted with the analytic formula
$R(\beta)=a+\frac{b}{e^{(\beta-c)/d}+1},$ (5)
where $\beta=k_{T}/m_{T}$ is pair velocity, and $a$, $b$, $c$, $d$ are free parameters of the fit. As particle velocity decreases, the radii seem to reach a saturated value, similar in a very broad range of $\beta$. Therefore such value could be used to determine an extrapolated value of the radii for ultra-soft pions, instead of the historically used dependence based on power-law $m_{T}$ scaling.
@@ V SUMMARY
In this paper, the model study of femtoscopic radii obtained from simulated data of iHKM and LQTH were shown. The results point out that the power-law character of the femtoscopic $m_{T}$-scaling is no longer valid for 1D studies in the low $m_{T}$ region (below 250 MeV$/c$). In the case of 3D studies, the radii are better described by the power-law dependence than 1D radii, however, the final conclusions are model-dependent. The iHKM model follows a power-law description in all 3 directions. On the other hand, for the LQTH model power-law description holds only in the longitudinal direction. Both models show that in low-$m_{T}$ region radii distribution tends to be more linear than power-law or even flat in some cases. This results in smaller radii than expected from the extrapolation of the power-law scaling from intermediate and high $m_{\rm{T}}$.
@@ VI ACKNOWLEDGMENTS
This work is funded by the Research University – Excellence Initiative of Warsaw University of Technology via the strategic funds of the Priority Research Centre of High Energy Physics and Experimental Techniques, the IDUB POSTDOC programme, the IDUB YOUNG-PW programme, Scientific Council of the discipline grant programme and by the Polish National Science Centre under agreements no. 2022/45/B/ST2/02029, and no. 2023/49/N/ST2/03525.

### v3

# Femtoscopy analysis of ultrasoft pion trap at energies available at the CERN Large Hadron Collider

@@ I INTRODUCTION
In experimental studies, one clearly observes certain types of universal scaling behaviour for the measured femtoscopy scales. For a given colliding system and collision energy a linear scaling of the femtoscopic radii is universally observed versus cube root of the final state mean particle multiplicity $\langle dN_{ch}/d\eta\rangle^{1/3}$ [16, 17, 18, 19, 20, 21, 22]. A similar scaling across collision energies and colliding systems is only approximate [23, 24]. The radii versus the pair transverse momentum exhibit a power-law-like scaling in pair transverse mass $m_{T}$ ($m_{T}$ = $\sqrt{k_{T}^{2}+m^{2}}$, where $k_{T}$ is pair mean transverse momentum, $k_{T}=|\textbf{p}_{T1}+\textbf{p}_{T2}|$/2, and $m$ is particle mass) [25, 26, 27, 28, 29, 16, 17, 30]. In heavy-ion collisions both these scalings are predicted by hydrodynamic models [31, 32]. Specifically, the $m_{T}$ scaling is explained as a direct consequence and one of the main signatures of the collective radial flow of the system.
@@
Thus, in this paper, we focus on studying the pion femtoscopic radii dependence on pair $m_{T}$, including the region of very low transverse masses, in Pb-Pb collisions at the LHC energy $\sqrt{s_{NN}}$ = 5.02 TeV simulated within the integrated HydroKinetic Model (iHKM) [31, 39, 40] and the hydrodynamics model [41] coupled to the statistical hadronisation code THERMINATOR2 (LHYQUID + THERMINATOR2) [42] — LQTH.
@@ II MODELS’ DESCRIPTION
The LQTH is similarly properly calibrated to describe the $\sqrt{s_{NN}}=5.02$ TeV Pb-Pb collisions at the LHC. The model assumes a second order viscous (3+1)D hydrodynamics evolution of the created fireball including shear and bulk viscosity as well as the Israel-Stewart stress corrections. The hydrodynamic expansion starts at 0.6 fm/c and lasts until the single (both chemical and kinetic) freeze-out at a selected temperature ($T_{p}\approx 140$ MeV), where statistical hadronization takes place. The system’s chemical composition and momentum spectra do not change anymore, except for the resonance propagation and decays, which are simulated by the THERMINATOR2 package. At the freeze-out temperature, particles are generated according to statistical rules from the freeze-out hypersurface following the Cooper-Frye formula. The model reproduces the heavy-ion collisions results including flow and femtoscopic measurements  [41, 47, 48]. All details about the model implementation and initial conditions can be found in  [41].
@@ III CONSTRUCTION OF THE CORRELATION FUNCTION
In experimental analysis, the procedure of obtaining the correlation function is based on making pairs of detected particles. One uses particles coming from the same event to build the correlated distribution (numerator), while for the background distribution (denominator) pairs are created using particles coming from different events (meaning that they retain all aspects of experimental acceptance, while not being correlated due to the mutual interaction).
@@ IV PION FEMTOSCOPY SCALES
[…]
@@
Figure 4: First-order derivative of the charged identical pions femtoscopic radii $m_{T}$ dependency, $dR/dm_{T}$, in the HKM (left) and the LQTH (right) models, for the three collision centralities (from top to bottom: $0-5\%,10-20\%,30-40\%$). Each panel contains derivatives of the three radii from the 3D study ($R_{\mathrm{out}}$ — blue, $R_{\mathrm{side}}$ — red, $R_{\mathrm{long}}$ — orange) and $R_{\mathrm{inv}}$ from the 1D study (green). Markers represent results obtained from simulations and the dashed lines are obtained for the expected power-low trend based on fits to simulation points above 250 MeV/$c$.
@@
Figure 5: The pion last interaction time and transverse coordinate in out direction of LCMS system for iHKM (left) and LQTH (right) simulations of Pb-Pb $\sqrt{s_{\rm{NN}}}$=5.02 TeV collisions. The upper panels show distributions for the lowest $k_{T}$ interval, the bottom ones — for the middle $k_{T}$ interval considered in this study. Sub-figures show: projection on $x_{out}$ direction (left), projection on time axis for the two $x_{out}$ ranges (center and right).
To better investigate the character of the radii $m_{\rm{T}}$ dependencies, they were differentiated, and the obtained derivatives are presented in Fig. 4 (for the three centrality classes). The derivative has been extracted in the range ±$\pm 0.07$ GeV/$c$ around each point in the figure. Most of the functions show similar behaviour, with a clear change in the derivative form at certain momenta. Namely, the derivative behaviour at low momenta ($m_{T}<0.25$ GeV/$c$), especially for the $R_{\rm{out}}$ and $R_{\rm{inv}}$ radii in LQTH model, is noticeably different from that of a power-law function derivative, typical for the higher momentum region and approximately followed even at low $m_{T}$ by the $dR_{\rm{long}}/dm_{T}$ dependencies. In more detail, in the iHKM one observes that the radius derivatives for the 1D case go flatter and are smaller by the absolute value than the $R_{\rm{long}}$ derivatives for all the collision centralities, while the out and the side curves go either slightly higher or slightly lower than the long one, depending on the collision type. In LQTH model we see a somewhat different situation, however, rather similar for all the centrality classes: all the derivatives, except for the long one, at low $m_{T}$ have much smaller absolute values than the power-law function and demonstrate nearly flat behaviour with small falls and rises, so that the radii dependency in this region can be sufficiently well approximated by a linear function.
@@
Note, that we see the mentioned effects of non-power-law transverse radii behaviour in the ultra-soft momentum region in “hydro plus hadronic cascade” models, where the dynamics is built in quasi-classical approximation. The Bose-Einstein quantum statistics effect is taken into account in the very final stage only, but even this way, using the femtoscopy method, one can see specific (delayed freeze-out) features of emission in the ultra-low-momentum region (see also Fig. 1 from [52]). In particular, as we already mentioned, the femtoscopy analysis brings us the signal about formation of a long-lived ultra-soft pion trap in the central part of the system created in ultra-relativistic heavy ion collision. To further check this signal, we built the distributions of the transverse out coordinate of pion last collision points vs its time in the LCMS system, see Fig. 5. It can be easily noticed that in both models the origin of particles forming low and high $k_{T}$ pairs is qualitatively different. Low $k_{T}$ pairs are mainly created of particles with low momenta most intensively emitted from the decaying system close to its center. To see this, one can compare the red part of the $x_{\rm{out}}$ distributions shown in the panels of Fig. 5 for low (top panels) and high (bottom panels) particle momenta, respectively. The red parts correspond to the emission coming from the hadronizing hydrodynamic tube itself, while the blue parts correspond to the emission at larger distances from the expanding hadron-resonance gas. The blue parts are rather similar at low and high momenta, while the red ones have a maximum in the center of the system for low momenta and at the periphery of the system for higher momenta. The red time distributions at low momenta have two maxima and larger mean emission time than those at high momenta having only one maximum near the system’s particlization time. The blue time distributions are similar for the low and high momentum cases.
The dependencies of the radii on $m_{T}$, together with the power-law fits, are presented for the three centrality classes in Fig. 6. The fitting was performed outside the “non-power-law” region observed in Fig. 4 (above 0.25 GeV/$c$). As one can see from Fig. 6, the power-law function in the form $am_{\rm{T}}^{-b}$ works well for all the radii at not very low momenta, and in case of long direction it describes even the ultra-soft region below 0.25 GeV/$c$. Other radii, especially the 1D ones, at very low momenta go below (in the iHKM sometimes above) the power-law fitting curve.
@@
Finally, we analyzed the femto-radii dependence on pair velocity, see Fig. 8 for different centrality intervals and for both used model approaches. In all cases, we notice a slow radii falling at low velocities. All the distributions were fitted with the analytic formula derived on purpose of this study
$R(\beta)=\frac{a}{e^{(\beta-b)/c}+1},$ (5)
where $\beta=k_{T}/m_{T}$ is pair velocity, and $a$, $b$, $c$ are free parameters of the fit. As particle velocity decreases, the radii seem to reach a saturated value, similar in a very broad range of $\beta$. Therefore such value could be used to determine an extrapolated value of the radii for ultra-soft pions, instead of the historically used dependence based on power-law $m_{T}$ scaling. It is particularly important for non-identical femtoscopy studies, where one considers pairs of particles with different masses moving at the same velocity – thus, having different momenta.
@@ V SUMMARY
In this paper, the model study of femtoscopic radii obtained from simulated data of iHKM and LQTH were shown. The results point out that the power-law character of the femtoscopic $m_{T}$-scaling is no longer valid for 1D studies in the low $m_{T}$ region (below 250 MeV$/c$). In the case of 3D studies, the radii are better described by the power-law dependence than 1D radii, however, the final conclusions are model-dependent. The iHKM model results are closer to a power-law behaviour, especially in the 3D analysis case. On the other hand, for the LQTH model one sees noticeable deviations from the power-law description everywhere, except for the longitudinal direction. Both models show that in low-$m_{\rm{T}}$ region radii distribution tends to be more linear than power-law or even flat in some cases. This results in smaller radii than expected from the extrapolation of the power-law scaling from intermediate and high $m_{\rm{T}}$.
@@ VI ACKNOWLEDGMENTS
This work is funded by the Research University – Excellence Initiative of Warsaw University of Technology via the strategic funds of the Priority Research Centre of High Energy Physics and Experimental Techniques, the IDUB POSTDOC programme, the IDUB YOUNG-PW programme, Scientific Council of the discipline grant programme and by the Polish National Science Centre under agreements no. 2022/45/B/ST2/02029, and no. 2023/49/N/ST2/03525. The work was also supported by a grant from the Simons Foundation (Grant Number 1290596, Yu.S. and V.S.).